7. Find 2 consecutive odd integers such that the sum of their squares is 130.
1
Expert's answer
2011-10-06T12:14:48-0400
7 and 9 fit the task because
7*7 + 9*9 = 130
(7*7+9*9=49+81=130)
and they are consecutive odd integers